the residents of a city voted on whether to raise property taxes. The ratio of yes to no votes was 5 to 8. If there were 6776 no votes, what was the total number of votes?
Answer by stanbon(57357)

the residents of a city voted on whether to raise property taxes. The ratio of yes to no votes was 5 to 8. If there were 6776 no votes, what was the total number of votes?
-------
5 to 8 is the same as 5x to 8x where 8x is the number of no votes.
----
If 8x = 6776, then x = 847
----
Total votes = 5x+8x = 13x
Tv = 13*847 = 11011
====================
Cheers,
Stan H.
====================